OpenStudy (anonymous):

Solve for x: 3/x + 1/2 = 5/6

OpenStudy (nurali):

\[\frac{ 3 }{ x }+\frac{ 1 }{ 2 }=\frac{ 5 }{ 6 }\] \[\frac{ 3 }{ x }=\frac{ 5 }{ 6 }-\frac{ 1 }{ 2 }\] \[\frac{ 3 }{ x }=\frac{ 1 }{ 3 }\] Cross Multiply \[x=3*3\] \[x=9\]

OpenStudy (anonymous):

How'd you get 1/3?

OpenStudy (anonymous):

@Nurali

OpenStudy (pradius):

pass 1/2 to the other side of the equal sign.. so u have 3/x alone

OpenStudy (pradius):

or better said .. subtract 1/2 in both sides

OpenStudy (nurali):

Subtract both side by 1/2

OpenStudy (anonymous):

Alright thanks!

OpenStudy (nurali):

My Pleasure.